ULTRAFAST STUDY OF BROMINE RADICAL IN SOLUTION: THE ROLE OF COMPLEXES AND VI-BRATIONAL EXCITATION

STACEY L. CARRIER, \underline{THOMAS J. PRESTON}, ANDREW C. CROWTHER, F. FLEMING CRIM, The University of Wisconsin - Madison Department of Chemistry, 1101 University Avenue, Madison, WI 53705.

The bromine radical is an interesting species to interrogate with vibrationally mediated chemistry. It forms a longlived complex in the liquid phase, and hydrogen abstraction from a solvent molecule is endothermic. Photolysis of a bromine precursor forms an iso-compound within 1 ps, which decays to a radical-solvent complex on a ∼10 ps timescale. We use these features to examine the reaction of the Br radical complex with an excited C-H oscillator. A potential reaction pathway that becomes available upon vibrational excitation is H abstraction from the excited oscillator by Br. Vibrational excitation can also cause dissociation of the Br complex. We follow the progress of these reactions using transient broadband UV-Vis spectroscopy.